# Precomputed Dominant Resource Fairness

### v1

@@ Abstract
Keywords Resource Distribution  $\cdot$ Dominant Resource Fairness  $\cdot$ Precomputed
@@ 1 Introduction
Dominant Resource Fairness (DRF) is a resource allocation scheme, which aims to provide fairness of allocation among users with heterogeneous demands, in a time-sharing system with multiple resource types, such as computing clouds (Ghodsi et al., 2011). The services offered by these systems may include processing power, memory, and bandwidth, all of which are expected to be distributed fairly among the clients, according to some definition of fairness.
Although the problem has been well-studied for single resource types (Keshav, 1997; Bertsekas et al., 1992), prior to DRF, less attention was paid to the multi-resource scenarios. The general approach was to adopt well established solutions for single resource allocation, such as Max-min Fairness (MF) and its derivatives, for a fixed bundle with differing amounts of resources from each resource type; e.g. standardised virtual machines. But this approach leads to inefficiencies, since the needs of different users are heterogeneous, where one needs predominantly one resource type, the other needs another. In the end, users ended up with unused resources that come in the bundle with the resources they need, where at the same time there were other users in the system in demand of those unused resources.
@@
DRF has acquired wide adoption in the distributed systems and cluster computation because of its properties, such as being Pareto optimal, Envy Free, Strategy Proof, and creating Sharing Incentive. In the present paper, we analyse the DRF algorithm as introduced by Ghodsi et al. (2011) in closer view, and based on the observations gathered thereby, introduce a new algorithm to approximate the DRF allocation in fewer steps. We call this new algorithm Precomputed Dominant Resource Fairness (PDRF), after its main working principle of precomputing the main allocation loop of DRF, and assigning resources to users without going through the tedious iterations of the loop.
One study that also needs to be mentioned at the onset, is that of Parkes et al. (2015), the method and the findings of which closely resemble to that of the present study, with some key differences that allow us to implement our newly proposed algorithm. The analysis method and the algorithm they propose, the Extended Dominant Resource Fairness (EDRF) are similarly based on precomputing the final allocation of DRF, but they modeled the underlying linear programming problem, rather than the algorithm itself. They construct a theoretical framework that allows them to provide more rigorous proofs of DRF properties, which is the main objective of the article. We elaborate on this in Section 3.2.
The rest of this paper is organised as follows: In the next section, we briefly review the literature for related work. In Section 3.4, we present our model, after reviewing DRF and EDRF in subsections 3.1 and 3.2, respectively, and introducing an analysis method in subsection 3.3. In section 4, we present the results of the tests we carried out with PDRF. In Section 5 we analyse PDRF performance. Section 6 consists of three subsections, in which we discuss higher order cycles in PDRF (6.1), Pareto Efficiency property (6.2), and weighting policies (6.3), which we believe, are important points to focus on critically. Finally, we conclude in Section 7.
@@ 2 Related Work
DRF is first proposed by Ghodsi et al. (2011), where they introduce the new notion of establishing an MF allocation among dominant shares, rather than a fixed resource bundle of different resource types. They also provide an algorithm to carry out the distribution, and produce empirical data to demonstrate the performance of the algorithm. In their theoretical sections, they compare the virtues of the newly introduced allocation scheme to the ones existing in the economy literature, for this is as much a problem in the field of economy as it is in the field of computer science. Their results clearly demonstrate that DRF possesses various advantages over its alternatives.
The model is challenged by Dolev et al. (2012). They developed an alternative fairness notion, which they refer to as no justified complaints, based on the assumption that the fairness of distribution of resources should be assessed from a global system point of view, considering the bottleneck property of resources. A resource is a bottleneck resource, if the availability of this resource is limited and there is a race condition for accessing it. They developed the Bottleneck Based Fairness (BBF) algorithm to implement this paradigm as an alternative to DRF, with relative advantages of each algorithm in different use cases.
Not only challenged, but DRF has also been adopted, generalised, contextualised, or extended to different cases. For example, Gutman and Nisan (2012), introduced an economical framework to compare and contrast the two systems described above. While doing so they also critically assessed both approaches from an economic point of view. Joe-Wong et al. (2013) investigates the trade-off between fairness and efficiency, providing a unifying framework with two families of fairness functions, DRF being a special case in one of them. Kash et al. (2014) takes a scenario in which the complete information on demands is not available, since users arrive and leave the system, and develops a Dynamic DRF (DDRF) to address the problem. The DDRF algorithm is improved for its performance in Li et al. (2015).
Finally, DRF has been reformulated by Parkes et al. (2015) to EDRF, in order to study its properties more rigorously. EDRF is also a solution to a special case, where the resources, as well as user demands are well-divisible, i.e. fractions of demands can be allocated to users, but for obvious reasons this is not widely applicable to real life scenarios, where half a task does not provide any added utility to the users. In the same article, they also introduce Sequential Min-max Algorithm to address the indivisible demands scenario. One more contribution of EDRF is that it addresses the situations where there are users who submit zero demands for one or more resources.
@@ 3.1 Dominant Resource Fairness
The DRF allocation scheme is build on the concept of dominant shares. The dominant share for each user is defined as the maximum of the fraction of her demands for different resource types to their respective reserves in the system. As exemplified in Ghodsi et al. (2011), in a system with $9$ CPUs and $18$ GB memory, the dominant share, $ds$ for user A, with the demand vector $\left\langle 1,4\right\rangle$ is memory, since the fractional demands, $fd$, of the user are $1/9$ and $4/18=2/9$, respectively. Similarly, the $ds$ of a user with the demand vector $\left\langle 3,1\right\rangle$ is for CPU, since $fd$’s are $3/9$ and $1/18$, respectively. The question then is, what is a fair allocation between the users, whose demands for the resources are defined as such.
If the number of tasks allocated to users A and B is denoted by $x$ and $y$, respectively, the DRF allocation is defined as the solution to the optimising problem below (Ghodsi et al., 2011, p. 4):
@@
$x+3y$$\leq$ 99 99 CPU Constraint
$4x+y$$\leq$ 1818 1818 Memory Constraint
@@
Designed to solve this problem, DRF is a progressive filling algorithm, applying MF on the users’ $ds$’s by looping on the total allocated $ds$’s of users, taking the least allocated at each iteration, and allocating one unit of her demand vector, assumed to correspond to one task, until one of the resources is exhausted. Although not explicitly stated, in addition to this mechanism, DRF utilises a tie-breaking policy to take the user with the higher $ds$, in instances when more than one user is allocated the same minimum share. This we see at the example given in the article, in the allocation process for the above mentioned case.
In the initial state, where both user’s allocated $ds$ is 00, the algorithm starts assigning one task to user B, which updates the state to 00 allocated $ds$ for A, and $1/3$ for B. Since now the least allocated $ds$ belongs to A, the algorithm picks her and assigns one task, updating the state to $2/9$ w.r.t. $1/3=3/9$. Still A has the lower allocated $ds$, so one more task is allocated to her, updating to $4/9$ w.r.t. $3/9$, and pushing B to the position of the user with the least allocated $ds$. Thusly the algorithm proceeds with B, and continuing further in the similar fashion, assigns $2/3$ dominant shares to each user, at which point the CPU is saturated and the algorithm stops.
As seen in the example above, DRF is a progressive filling algorithm, since it aims to keep the growth of allocation among different agents at the same pace. DRF reduces to MF, and also to BBF (Dolev et al., 2012), in cases where all users’ dominant resources are the same.
@@ 3.2 Extended Dominant Resource Fairness
The study by Parkes et al. (2015) introduces an alternative view to DRF. Instead of an iterative algorithm, they aim to precompute the resulting distribution, with a linear program. For doing so, they employ two notions in the setting. First, in a given user demand vector, the ratio between the entries for different resources is fixed. Second, the operating principles of progressive filling algorithm dictates that the pace at which the allocation for dominant shares of different users grow should be equated. Departing from these notions, it is possible to normalise each demand vector to its dominant share, and abstract that they will all be allocated simultaneously. In other words, if we scale all the dominant shares to $1$, scale all other fractional shares accordingly, we end up in a setting in which at each step, every user takes $1$ normalised allocation, and fairness among the dominant shares is kept intact.
The question then is, what is the maximum number of such allocations that can be made before depleting one of the resources? The answer starts with normalising the vector of fractional demands,$fd_{ir}$ to the dominant share, $ds_{i}$, which is denoted $d_{ir}$ and defined more explicitly as:
@@ 3.3 Cyclic Structure of the DRF Main Loop
where $\lvert l\rvert$ denotes the length of the cycle, $\operatorname{lcm}DS$ denotes the least common multiple of the set of all dominant shares, $DS$, and $ds_{i}$ is the dominant share of user $i$. The individual occurrences of each demand within the cycle is given by the $i$-$th$ term of the summation for user $i$. In the cases where all other dominant shares are integer fractions of $ds^{*}$, like in the example above, since $lcmS$ is equal to $ds^{*}$, the latter part is equivalent to the ratio of each $ds_{i}$ to $ds^{*}$, and the length of the cycle is equal to the sum of all these ratios, plus $1$, for the single occurrence of $ds^{*}$.
@@
where $\lvert l^{*}\rvert$ denotes the length of, what we would like to call, the basic subcycle. The other subcycles are characterised by including an additional occurrence of one or more of the lower dominant share demands, the frequency of which is determined by the decimal part of the ratio of $ds^{*}$ to each dominant share, i.e., the $i$-$th$ term of the expression above for demand $i$. Trivially, the first subcycle will always have an extra occurrence, if the ratio has a decimal part.
@@
3.4 Precomputed Dominant Resource Fairness
@@
$k=\min_{r\in R}\left\{\frac{r}{\sum_{i}\left(\frac{ds^{*}}{ds_{i}}\cdot d_{ir}\right)}\right\}=\min_{r\in R}\left\{\frac{r}{ds^{*}\cdot\sum_{i}\frac{d_{ir}}{ds_{i}}}\right\}=\frac{1}{ds^{*}}\cdot\min_{r\in R}\left\{\frac{r}{\sum_{i}\frac{d_{ir}}{ds_{i}}}\right\}$
@@
but presumably for numerical reasons, such a simplification leads to less accurate results, and surprisingly, poorer algorithm performance. Thus we do not implement this simplification.
Algorithm 1 Precomputed DRF
1:$R=\left\langle r_{1},r_{2},r_{3},\dots,r_{m}\right\rangle$triangleright resource reserves
2:$C=\left\langle c_{1},c_{2},c_{3},\dots,c_{m}\right\rangle$triangleright consumed amounts of resources
3:$S=\left\langle s_{1},s_{2},s_{3},\dots,s_{n}\right\rangle$triangleright set of dominant shares
4:$U_{i}=\left\langle u_{i1},u_{i2},u_{i3},\dots,u_{im}\right\rangle$triangleright allocation vector for the user i
5:$D_{i}=\left\langle d_{i1},d_{i2},d_{i3},\dots,d_{im}\right\rangle$triangleright demand vector for the user i
@@
7:pick user with the highest dominant share $\text{{pick} user with the highest dominant share }ds^{*}$
@@
9:     $d_{total}\leftarrow d_{total}+\frac{ds^{*}}{ds_{i}}\cdot D_{i}$triangleright Cumulate the total size of demand in a subloop
@@
12:$k=\min_{r}\left\{\frac{r}{d_{total}}\right\}$triangleright find maximum available iterations of the subloop
@@
15:     $U_{i}\leftarrow U_{i}+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$triangleright allocate each user accordingly, for a given subloop
16:     $C\leftarrow C+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$triangleright update the consumed resources vector
@@
The pseudocode of PDRF is presented in Algorithm 1. A sample implementation can be accessed at this github repository.
@@ 4 Results
In this section we present the results of the tests we have carried out by samples drawn from the discrete uniform distribution. $1000$ users are assigned random demands, and resource reserves are likewise assigned randomly. We tested for various numbers of resource types, demand intervals and resource reserve intervals. Here we are presenting a sample which we consider representative. We run each test $1000$ times, and collected the averages, standard deviations, and maximums, presented in Table 1 and Table 2. The scripts that we used to run the tests can be accessed at the same repository, here.
The results show that the deviations of the PDRF allocation from the DRF allocation are limited to a small number of tasks per user, predominantly by $1$, and very rarely exceeding $2$. However, the error is two tailed, meaning, the deviations may either be underallocations or overallocations, the latter of which is a rather serious drawback, since, unlike underallocations, they cannot be compensated for with a final iteration of DRF. Nevertheless, the rate at which these occur are negligibly low.
In the tests carried out with resource intervals drawn from $[50,000-100,000]$ interval and demands are drawn from $[1-10]$ interval, on avarage, $47.7\%$ ($\mu=477.623,n=1000$) of the users are underallocated $1$ task, with a standard deviation of $\sigma=86,373$. The figures for demand intervals $[1-20]$ and $[10-20]$ are similarly, $\mu=474.932,\sigma=38.381$ and $\mu=486.521,\sigma=50.526$, respectively. The overallocations are for demand intervals $[1-10]$ and $[1-20]$ are $\mu=0.694,\sigma=2.137$ and $\mu=0.192,\sigma=0.526$, respectively.
1,000 Users, 10 Resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Underallocations Overallocations
Demand Intervals 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev. 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev.
1 – 10 477.623 0 0 1 477.623 86.373 0.693 0.001 0 2 0.694 2.137
1 – 20 474.932 0 0 1 474.932 38.381 0.191 0.001 0 2 0.192 0.526
10 – 20 486.521 0 0 1 486.521 50.526 0 0 0 0 0 0
Table 1: Average Case Scenarios
@@
In more general terms, for a given demand, as the ratio of maximum dominant share to its dominant share increases, the number of overallocations increase proportionally. This is caused by large demand and resource reserve intervals which lead to increased variances in the distribution of dominant shares, and mitigated with the increasing number of resource types, since it leads to lower variance among dominant shares. The worst configuration we could cook were with $2$ resource types, as seen in Table 2.
The reason for the algorithm’s implicit favouring of the smaller dominant shares is due to the the rather abstract notion of partial cycles, which are reflected in the decimal points of the number of cycle iterations ($k$). If the exact ordering of the cycle were available, those numbers would indicate the fraction of the cycle that would execute from the beginning, e.g. a decimal point of 0.5 would indicate, first half of the allocations would be done and the cycle would terminate. In the absence of the exact ordering of allocations, the users exploit the partial cycles inversely proportional to the ratio of the $ds^{*}$ to their $ds$.
1000 Users, 2 Resources, Resource Reserves Drawn From [50000 – 100000] Interval
Underallocations Overallocations
Demand Intervals 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev. 1 Task 2 Tasks Tasks > 2 Maximum Avg. St. Dev.
1-100 423.517 0 0 1 423.517 36.429 44.689 5.588 0.48 3 50.407 25.814
1-1000 407.902 0 0 1 407.902 26.373 56.541 14.519 12.989 32 83.959 36.376
Table 2: Worse Case Scenarios

### v2

@@ Abstract
keywords:
Precomputed , Dominant Resource Fairness , Resource Allocation , Analysis.
journal: arxiv.org
@@ 1 Introduction
Dominant Resource Fairness (DRF) is a resource allocation scheme, which aims to provide fairness of allocation among users with heterogeneous demands, in a time-sharing system with multiple resource types, such as computing clouds [1]. The services offered by these systems may include processing power, memory, and bandwidth, all of which are expected to be distributed fairly among the clients, according to some definition of fairness.
Although the problem has been well-studied for single resource types [2, 3], prior to DRF, less attention was paid to the multi-resource scenarios. The general approach was to adopt well established solutions for single resource allocation, such as Max-min Fairness (MF) and its derivatives, for a fixed bundle with differing amounts of resources from each resource type; e.g. standardised virtual machines. But this approach leads to inefficiencies, since the needs of different users are heterogeneous, where one needs predominantly one resource type, the other needs another. In the end, users ended up with unused resources that come in the bundle with the resources they need, where at the same time there were other users in the system in demand of those unused resources.
@@
DRF has acquired wide adoption in the distributed systems and cluster computation because of its properties, such as being Pareto Efficient, Envy Free, Strategy Proof, and creating Sharing Incentive.
In the present paper, we analyse the DRF algorithm as introduced in [1] in closer view, and based on the observations gathered thereby, introduce a new algorithm to approximate the DRF allocation in fewer steps. We call this new algorithm Precomputed Dominant Resource Fairness (PDRF), after its main working principle of precomputing the main allocation loop of DRF, and assigning resources to users without going through the tedious iterations of the loop.
One study that also needs to be mentioned at the onset, is [4], the method and the findings of which closely resemble to that of the present study, with some key differences that allow us to implement our newly proposed algorithm. The analysis method and the algorithm they propose, the Extended Dominant Resource Fairness (EDRF) are similarly based on precomputing the final allocation of DRF, but they modeled the underlying linear programming problem, rather than the algorithm itself. They construct a theoretical framework that allows them to provide more rigorous proofs of DRF properties, which is the main objective of the article. We elaborate on this in Section 3.2.
The rest of this paper is organised as follows: In the next section, we briefly review the literature for related work. In Section 4, we present our model, after reviewing DRF and EDRF in subsections 3.1 and 3.2, respectively, and introducing an analysis method in subsection 3.3. We introduce PDRF in Section 4. In section 5, we present the results of the tests we carried out with PDRF, and departing on the observations gathered by there, in Section 6, we present an heuristic method to improve pure PDRF results. In Section 7 we analyse PDRF performance. Section 8 consists of three subsections, in which we discuss higher order cycles in PDRF (8.1), Pareto Efficiency property (8.2), and weighting policies (8.3), which we believe, are important points to focus on critically. Finally, we conclude in Section 9.
@@ 2 Related Work
DRF is first proposed in [1], where they introduce the new notion of establishing an MF allocation among dominant shares, rather than a fixed resource bundle of different resource types. They also provide an algorithm to carry out the distribution, and produce empirical data to demonstrate the performance of the algorithm. In their theoretical sections, they compare the virtues of the newly introduced allocation scheme to the ones existing in the economy literature, for this is as much a problem in the field of economy as it is in the field of computer science. Their results clearly demonstrate that DRF possesses various advantages over its alternatives.
The model is challenged in [5]. They developed an alternative fairness notion, which they refer to as no justified complaints, based on the assumption that the fairness of distribution of resources should be assessed from a global system point of view, considering the bottleneck property of resources. A resource is a bottleneck resource, if the availability of this resource is limited and there is a race condition for accessing it. They developed the Bottleneck Based Fairness (BBF) algorithm to implement this paradigm as an alternative to DRF, with relative advantages of each algorithm in different use cases.
Not only challenged, but DRF has also been adopted, generalised, contextualised, or extended to different cases. For example, Gutman and Nisan [6], introduced an economical framework to compare and contrast the two systems described above. While doing so they also critically assessed both approaches from an economic point of view. Joe-Wong and colleagues [7] investigates the trade-off between fairness and efficiency, providing a unifying framework with two families of fairness functions, DRF being a special case in one of them. Kash [8] takes a scenario in which the complete information on demands is not available, since users arrive and leave the system, and develops a Dynamic DRF (DDRF) to address the problem. The DDRF algorithm is improved for its performance in [9].
Finally, DRF has been reformulated by Parkes and colleagues [4] to EDRF, in order to study its properties more rigorously. EDRF is also a solution to a special case, where the resources, as well as user demands are well-divisible, i.e. fractions of demands can be allocated to users, but for obvious reasons this is not widely applicable to real life scenarios, where half a task does not provide any added utility to the users. In the same article, they also introduce Sequential Min-max Algorithm to address the indivisible demands scenario. One more contribution of EDRF is that it addresses the situations where there are users who submit zero demands for one or more resources.
@@ 3.1 Dominant Resource Fairness
The DRF allocation scheme is build on the concept of dominant shares. The dominant share for each user is defined as the maximum of the fraction of her demands for different resource types to their respective reserves in the system. As exemplified in [1], in a system with $9$ CPUs and $18$ GB memory, the dominant share, $ds$ for user A, with the demand vector $\left\langle 1,4\right\rangle$ is memory, since the fractional demands, $fd$, of the user are $1/9$ and $4/18=2/9$, respectively. Similarly, the $ds$ of a user with the demand vector $\left\langle 3,1\right\rangle$ is for CPU, since $fd$’s are $3/9$ and $1/18$, respectively. The question then is, what is a fair allocation between the users, whose demands for the resources are defined as such.
If the number of tasks allocated to users A and B is denoted by $x$ and $y$, respectively, the DRF allocation is defined as the solution to the optimising problem below [1, p. 4]:
@@
$x+3y$ $\leq$ $9$ CPU Constraint
$4x+y$ $\leq$ $18$ Memory Constraint
@@
Designed to solve this problem, DRF applies MF principle to maximise the minimum share, on the users’ $ds$’s by looping on the total allocated $ds$’s of users, taking the least allocated at each iteration, and allocating one unit of her demand vector, assumed to correspond to one task, until one of the resources is exhausted.
Although not explicitly stated, in addition to this mechanism, DRF utilises a tie-breaking policy to take the user with the higher $ds$, in instances when more than one user is allocated the same minimum share. This we see at the example given in the article, in the allocation process for the above mentioned case.
In the initial state, where both user’s allocated $ds$ is 0, the algorithm starts assigning one task to user B, which updates the state to 0 allocated $ds$ for A, and $1/3$ for B. Since now the least allocated $ds$ belongs to A, the algorithm picks her and assigns one task, updating the state to $2/9$ w.r.t. $1/3=3/9$. Still A has the lower allocated $ds$, so one more task is allocated to her, updating to $4/9$ w.r.t. $3/9$, and pushing B to the position of the user with the least allocated $ds$. Thusly the algorithm proceeds with B, and continuing further in the similar fashion, assigns $2/3$ dominant shares to each user, at which point the CPU is saturated and the algorithm stops.
As seen in the example above, DRF adopts progressive filling strategy, by aiming to keep the growth of allocation among different agents at the same pace. DRF reduces to MF, and also to BBF [5], in cases where all users’ dominant resources are the same.
@@ 3.2 Extended Dominant Resource Fairness
The study by [4] introduces an alternative view of DRF. Instead of an iterative algorithm, they aim to precompute the resulting distribution, with a linear program. For doing so, they employ two notions in the setting. First, in a given user demand vector, the ratio between the entries for different resources is fixed. Second, the operating principles of progressive filling algorithm dictates that the pace at which the allocation for dominant shares of different users grow should be equated.
Departing from these notions, it is possible to normalise each demand vector to its dominant share, and abstract that they will all be allocated simultaneously. In other words, if we scale all the dominant shares to $1$, scale all other fractional shares accordingly, we end up in a setting in which at each step, every user takes $1$ normalised allocation, and fairness among the dominant shares is kept intact.
[…]
@@ 3.3 Cyclic Structure of the DRF Main Loop
where $\lvert l\rvert$ denotes the length of the cycle, $\operatorname{lcm}DS$ denotes the least common multiple of the set of all dominant shares, $DS$, and $ds_{i}$ is the dominant share of user $i$. The individual occurrences of each demand within the cycle is given by the $i$-thth term of the summation for user $i$. In the cases where all other dominant shares are integer fractions of $ds^{*}$, like in the example above, since $lcmS$ is equal to $ds^{*}$, the latter part is equivalent to the ratio of each $ds_{i}$ to $ds^{*}$, and the length of the cycle is equal to the sum of all these ratios, plus $1$, for the single occurrence of $ds^{*}$.
@@
where $\lvert l^{*}\rvert$ denotes the length of, what we would like to call, the basic subcycle. The other subcycles are characterised by including an additional occurrence of one or more of the lower dominant share demands, the frequency of which is determined by the decimal part of the ratio of $ds^{*}$ to each dominant share, i.e. the $i$-thth term of the expression above for demand $i$. Trivially, the first subcycle will always have an extra occurrence, if the ratio has a decimal part.
@@
4 Precomputed Dominant Resource Fairness
@@
$k=\min_{r\in R}\left\{\frac{r}{\sum_{i}\left(\frac{ds^{*}}{ds_{i}}\cdot d_{ir}\right)}\right\}=\frac{1}{ds^{*}}\cdot\min_{r\in R}\left\{\frac{r}{\sum_{i}\frac{d_{ir}}{ds_{i}}}\right\}$
@@
but for numerical reasons, such a simplification leads to less accurate results, and surprisingly, poorer algorithm performance. Thus we do not implement this simplification.
[…]
@@
[…]
@@
9:  $d_{t}\leftarrow d_{t}+\frac{ds^{*}}{ds_{i}}\cdot D_{i}$ $\triangleright$ demands total
@@
[…]
@@
15:  $U_{i}\leftarrow U_{i}+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$ $\triangleright$ allocation
16:  $C\leftarrow C+\left\lfloor k\cdot\frac{ds^{*}}{ds_{i}}\right\rfloor\cdot D_{i}$ $\triangleright$ update consumed
@@
Algorithm 1 Precomputed DRF
The pseudocode of PDRF is presented in Algorithm 1. A sample implementation can be accessed at this github repository [10].
@@ 5 Results
In this section we present the results of the tests we have carried out by samples drawn from the discrete uniform distribution.
$1000$ users are assigned random demands, and resource reserves are likewise assigned randomly. We tested for various numbers of resource types, demand intervals and resource reserve intervals. Here, we are presenting a sample which we consider representative. We run each test $1000$ times, and collected the averages, standard deviations, and maximums, presented in Tables 1 - 4. The scripts that we used to run the tests can be accessed at the same repository, here [10].
The results show that the deviations of the PDRF allocation from the DRF allocation are limited to a small number of tasks per user, predominantly by $1$, and very rarely exceeding $2$. However, the error is two tailed, meaning, the deviations may either be underallocations or overallocations, the latter of which is a rather serious drawback, since, unlike underallocations, they cannot be compensated with additional allocations. Nevertheless, the rate at which these occur are negligibly low.
In the tests carried out with resource intervals drawn from [50,000100,000][50,000-100,000] interval and demands are drawn from [110][1-10] interval, on avarage, $47.7\%$ ($\mu=477.623$, $n=1000$) of the users are underallocated $1$ task, with a standard deviation of $\sigma=86,373$. The figures for demand intervals [120][1-20] and [1020][10-20] are similarly, $\mu=474.932$, $\sigma=38.381$ and $\mu=486.521$, $\sigma=50.526$, respectively. The overallocations are for demand intervals [110][1-10] and [120][1-20] are $\mu=0.694$, $\sigma=2.137$ and $\mu=0.192$, $\sigma=0.526$, respectively.
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1 – 10 477.623 0 0 1 477.623 86.373
1 – 20 474.932 0 0 1 474.932 38.381
10 – 20 486.521 0 0 1 486.521 50.526
Table 1: Average Case Scenarios - Underallocations: Tests Done with 1000 users, 10 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1 – 10 0.693 0.001 0 2 0.694 2.137
1 – 20 0.191 0.001 0 2 0.192 0.526
10 – 20 0 0 0 0 0 0
Table 2: Average Case Scenarios - Overallocations: Tests Done with 1000 users, 10 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
@@
In more general terms, for a given demand, as the ratio of maximum dominant share to its dominant share increases, the number of overallocations increase proportionally. This is caused by large demand and resource reserve intervals which lead to increased variances in the distribution of dominant shares, and mitigated with the increasing number of resource types, since it leads to lower variance among dominant shares. The worst configuration we could cook were with $2$ resource types, as seen in Tables 3 and 4.
The reason for the algorithm’s implicit favouring of the lower dominant shares is due to the the rather abstract notion of partial cycles, which are reflected in the decimal points of the number of cycle iterations. If the exact ordering of the cycle were available, those numbers would indicate the fraction of the cycle that would execute from the beginning, e.g. a decimal point of 0.5 would indicate that the first half of the allocations would be done and the cycle would terminate. In the absence of the exact ordering of allocations, the users exploit the partial cycles inversely proportional to the ratio of the $ds^{*}$ to their $ds$.
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1-100 428.653 0 0 1 428.653 21.919
1-1000 430.86 0 0 1 430.86 37.002
Table 3: Worst Case Scenarios - Underallocations: Tests Done with 1000 users, 2 resources, Resource Reserves Drawn From [500,000 – 1,000,000] Interval
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1-100 40.267 10.12 9.239 35 59.626 10.790
1-1000 33.061 14.519 8.256 610 49.819 27.055
Table 4: Worst Case Scenarios - Overallocations: Tests Done with 1000 users, 2 resources, Resource Reserves Drawn From [500,000 – 1,000,000] Interval
6 An Heuristic Method for Allocating Remaining Resources
Based on the observations we made from the results, a simple heuristic approach to improve accuracy is to allocate $1$ task to each user, in the ascending order of allocated dominant shares, to as much users as possible, since the number of underallocations is invariably by $1$.
With such an extra step the rate of underallocation falls to $\mu=1.2\%$ and $\sigma=.559$, and overallocations rises to $\mu=1.3\%$ and $\sigma=.543$, on average. In the worst case, these numbers become $\mu=1.6\%$ and $\sigma=.257$, and $5.8\%$ and $\sigma=.638$ respectively. The results with the addition of the heuristic method may be seen in more detail in tables 5 - 8.
In addition to the improved accuracy, with the additional heuristic step, the algorithm also reaches Pareto Efficiency. Since overallocations are invariably for the smaller $ds$ values, PDRF makes higher numbers of task allocations with respect to DRF.
Nevertheless it should be noted that this heuristic method is specific to the discrete uniform distribution, and there is no guarantee it will work under different distributions.
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1 – 10 17.600 0 0 1 17.600 15.114
1 – 20 6.436 0 0 1 6.436 5.937
10 – 20 11.459 0 0 1 11.459 8.710
Table 5: Average Case Scenarios - Underallocations: Tests Done with 1000 users, 10 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1 – 10 19.116 0 0 1 19.116 15.266
1 – 20 8.521 0 0 1 8.521 6.160
10 – 20 11.716 0 0 1 11.716 8.764
Table 6: Average Case Scenarios - Overallocations: Tests Done with 1000 users, 10 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1-100 18.412 0 0 1 18.412 5.694
1-1000 14.773 0 0 1 14.773 9.928
Table 7: Worst Case Scenarios - Underallocations: Tests Done with 1000 users, 2 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval
Demand 1 task 2 tasks More Max. Avg. St. Dev.
1-100 43.608 9.71 9.53 34 62.848 11.139
1-1000 36.698 8.413 8.714 259 53.825 53.825
Table 8: Worst Case Scenarios - Overallocations: Tests Done with 1000 users, 2 resources, Resource Reserves Drawn From [50,000 – 100,000] Interval